### v1

@@ 3 Methods
Generalizing Volume Rendering Priors. We use the volume rendering prior represented by $\theta_{w}$ of $f_{w}$ to estimate a UDF $f_{u}$ from a set of RGB images $\{I_{j}\}_{j=1}^{J}$ of an unseen scene. We learn $f_{u}$ by minimizing the rendering errors on RGB images.
@@
Specifically, for a ray $V_{k}$, we hierarchically sample $N$=128 queries $\{q_{n}\}$ to render RGB values through volume rendering using Eq.(1). Similarly, we calculate the opaque density at each location $q_{n}$ using Eq.(2), but using unsigned distances predicted by $f_{u}$ as $\sigma_{n}=f_{w}(\{\delta_{m}\}_{m=1}^{M},\{f_{u}(q_{m}^{\prime})_{m=1}^{M}\})$, not the ground truth ones when learning $f_{w}$, and keeping the parameters $\theta_{w}$ of $f_{w}$ fixed during the generalizing procedure. Moreover, we use two neural networks to model the UDF $f_{u}$ and the color funtion $f_{c}$ parameterized by $\theta_{u}$ and $\theta_{c}$, respectively. We jointly learn $\theta_{u}$ and $\theta_{c}$ by minimizing the errors between rendered RGB images $\{\tilde{I}_{j}\}_{j=1}^{J}$ and the ground truth below,
$\mathcal{L}_{rgb}=\sum\nolimits_{j=1}^{J}||I_{j}-\tilde{I}_{j}||.$ (4)
@@
where $\lambda$ is a balance weight and set to 0.1 following previous work [47].
@@
Using the learned parameters $\theta_{u}$, we use the method introduced in [17] to extract the zero-level set of $f_{u}$ as the surface.
Implementation Details. We implement our volume rendering priors network $f_{w}$ as a 6-layer MLP with 256 hidden units and skip connections. Similar to previous work [47, 30], the UDF function $f_{u}$ is an 8-layer MLP with skip connections and the color function $f_{c}$ is a 4-layer MLP with 256 hidden units. To control the smoothness of the UDF learning, similar as the trainable variance in NeuS [47], we utilize two parameter sets of $f_{w}$ for early and later UDF inference stage, respectively. More implementation details can be found in the supplementary materials.
4 Experiments
We evaluate our method in surface reconstruction from multi-view RGB images. We report numerical and visual comparisons with the latest methods of learning UDFs under the same experimental setting. We also report ablation studies to justify the effectiveness of our modules and the effect of key parameters.
4.1 Experiment Settings
@@
Data for Learning Priors. We select one object from “car” category of ShapeNet dataset [5] and one from DeepFashion3D dataset [64] to form our training dataset for learning volume rendering priors. Note that there is no overlap between the selected object and the testing objects. Our ablation studies demonstrate that these two objects are sufficient to learn accurate volume rendering priors with good generalization capabilities across various shape categories. For each object, we first convert it into a normalized watertight mesh and then render 100 depth images with 600$\times$600 resolution from uniformly distributed camera viewpoints on a unit sphere. Without additional annotation, we utilize the volume rendering priors pre-trained on these two shapes to report our results.
Datasets for Evaluations. We evaluate our method on four datasets including DeepFashion3D (DF3D) [64], DTU [21], Replica [42] and real-captured datasets. For DF3D dataset, we follow [30] and use the same 12 garments from different categories. For DTU dataset, we use the same 15 scenes that are widely used by previous studies. And we use all the 8 scenes in Replica dataset. We also report results on real scans from NeUDF [29] and the ones shot by ourselves.
Baselines. We compare our method with the state-of-the-art methods which use different differentiable renderers to reconstruct open surfaces, including NeuralUDF [30], NeUDF [29] and NeAT [35]. We also report the results of NeuS [47] and COLMAP [41] as baselines. Note that NeuralUDF uses additional patch loss [31] to fine-tune the resulted meshes, which is not the primary contribution of the differentiable renderer or used by other methods. Hence, for fair comparison, we report the results of NeuralUDF without fine-tuning across all datasets. However, we still perform additional experiments in the supplementary to show that our method, when getting fine-tuned using the patch loss, outperforms NeuralUDF under the same experimental conditions.
@@
Table 1: Numerical comparisons in all ShapeNet categories.
@@
NeuS-UDF[47] 3.46$\pm$1.49 2.67$\pm$1.27 2.27$\pm$1.09
NeUDF[29] 1.67$\pm$0.31 4.19$\pm$2.02 0.89$\pm$0.14
NeuralUDF[30] 1.28$\pm$0.28 30.65$\pm$7.59 0.61$\pm$0.12
NeuS-SDF[47] 0.97$\pm$0.67 0.60$\pm$0.56 0.52$\pm$0.47
Ours 0.41$\pm$0.19 0.70$\pm$0.63 0.12$\pm$0.05
Metrics. For DTU dataset and DF3D dataset, we use Chamfer Distance (CD) as the metric. For Replica dataset, we report CD, Normal Consistency (N.C.) and F1-score following previous works [55, 3]. Moreover, we report the rendering errors in Tab. 1 using depth L1 distance, mask errors with cross entropy and L1 distance. The definitions of the metrics are provided in the supplementary materials.
Figure 6: Visual comparisons on open surface reconstructions with error maps on DeepFashion3D [64] dataset (NeAT uses additional mask supervision). The large reconstruction errors are shown in yellow on error maps.
Figure 7: Visual comparisons of error maps on DTU [21] dataset. The transition from blue to yellow indicates larger reconstruction errors.
Figure 8: Qualitative comparisons on Replica [42] dataset (NeAT uses additional mask supervision). Our method outperforms other methods on complex indoor scenes while other UDF-based methods struggle to recover complete and smooth surfaces.
4.2 Comparisons with the Latest Methods
Table 2: Quantitative evaluations on DF3D[64], DTU[21] and Replica[42] datasets. Note that NeAT uses mask supervision.
@@ 4.2 Comparisons with the Latest Methods
Ours 1.71 0.85 0.04 0.90 0.80
Results on ShapeNet. Tab. 1 reports numerical comparisons in the experiment in Fig. 2 in terms of 3 metrics. We report the averages and variances over all 55 shapes. We render depth images and mask images by forwarding ground truth unsigned distances at the same set of queries as other methods with the learned prior. We calculate the L1-error and cross entropy error between predicted images and ground truth ones. We achieve the best accuracy among all renderers for UDFs and SDFs.
@@
Results on DF3D. We first report evaluations on DF3D (CD ×$\times 10^{-3}$). Numerical comparison in Tab. 2 indicates that our learned prior produces the lowest CD errors among all handcrafted renderers. The visual comparison in Fig. 6 details our superiority on reconstruction with error maps. We see that our prior helps the network to recover not only smoother surfaces at most areas but also sharper edges at wrinkles. We also outperform NeAT [35] which uses additional mask supervisions to learn local SDFs and reconstruct open surfaces. Please see our supplementary materials for evaluations on each scene.
Figure 10: Visualization of our real-captured scenes.
Results on DTU. Tab. 2 reports our evaluation on DTU. Our reconstructions produce the lowest CD errors with our pre-trained prior. Although the shapes used to learn the prior are not related to any scenes in DTU, our prior comes from sets of queries in a local window on a ray, which are more general to unsigned distances along a ray in unobserved scenes. Hence, our prior produces excellent generalization ability. Visual comparisons in Fig. 7 detail our reconstruction accuracy. Please see our supplementary materials for evaluations on each scene.

### v2

@@ III-D Generalizing Volume Rendering Priors
We use the volume rendering prior represented by $\theta_{w}$ of $f_{w}$ to estimate a UDF $f_{u}$ from a set of RGB images $\{I_{j}\}_{j=1}^{J}$ of an unseen scene. We learn $f_{u}$ by minimizing the rendering errors on RGB images.
@@
Specifically, for a ray $V_{k}$, we hierarchically sample NN=128 queries $\{q_{n}\}$ to render RGB values through volume rendering using Eq. (1). Similarly, we calculate the opaque density at each location $q_{n}$ using Eq. (2), but using unsigned distances predicted by $f_{u}$ as $\sigma_{n}=f_{w}(\{\delta_{m}\}_{m=1}^{M},\{f_{u}(q_{m}^{\prime})_{m=1}^{M}\})$, not the ground truth ones when learning $f_{w}$, and keeping the parameters $\theta_{w}$ of $f_{w}$ fixed during the generalizing procedure. Although our model is trained on ground truth UDFs, it can effectively converge during generalization when initialized from a spherical UDF field. This is because our volume rendering prior learns a local mapping from UDFs to opaque densities, and is exposed to abundant local UDF patterns through multi-view optimization and random sampling during training. Under this setting, the multi-view photometric loss is back-propagated through volume rendering and $f_{w}$ to the entire UDF field, which drives the UDFs to gradually converge. Moreover, we use two neural networks to model the UDF $f_{u}$ and the color function $f_{c}$ parameterized by $\theta_{u}$ and $\theta_{c}$, respectively. We jointly learn $\theta_{u}$ and $\theta_{c}$ by minimizing the errors between rendered RGB images $\{\tilde{I}_{j}\}_{j=1}^{J}$ and the ground truth below,
[…]
@@
where $\lambda$ is a balance weight and set to 0.1 following previous work [62]. Using the learned parameters $\theta_{u}$, we use the method introduced in MeshUDF [21] to extract the zero-level set of $f_{u}$ as the surface.
To facilitate the smoothness learning of neural distance functions, existing methods typically introduce a learnable variance parameter to control the sharpness of the mapping functions [62, 68, 40]. However, there is no explicit supervision that defines the correspondence between the variance and the UDF mapping functions. To address this issue, we propose a two-stage parameter strategy that employs two parameter sets of $f_{w}$ for the early and late stages of UDF inference. Specifically, the parameters of $f_{w}$ are saved twice during training, once in the middle and once at the end. The weight decay of the optimizer is reduced as training progresses. We observe that in the middle stage with a large weight decay, the network $f_{w}$ exhibits an underfitting behavior and produces a loose mapping that is less sensitive to UDF variations, which enables it to perceive a wide range of potential surface locations, as shown in “stage 1” of Fig. 8. At the end of training with a small weight decay, the network learns sharp boundaries around the surfaces, resulting in rendering weights that are more concentrated on the zero-level set, as shown in “stage 2” of Fig. 8. During the UDF inference, we firstly load the first parameters set as our prior, which recovers a coarse UDF representing the overall geometry. Then we load the second parameters set, which will refine the UDFs near the already determined surfaces, producing fine details and sharp edges.
Figure 9: Illustration of our UDF upsampling strategy. We significantly mitigate the bias in hierarchical sampling by introducing an auxiliary point sampling prior to correct the sampling density in non-intersection regions.
Figure 10: Illustration of our uniform upsampling strategy. We group the upsampled points that fall within the same interval formed by the original sampled points, and re-order them to evenly distribute them within the interval.
III-E Hierarchical Volume Sampling for UDFs
Hierarchical volume sampling is an important strategy of improving rendering efficiency and 3D perception in volume rendering. NeuS [62] uses S-density function with fixed standard deviations for importance sampling, while the situation becomes more complex when it comes to UDF, as UDF which is unlike SDF, lacks an explicit distinction of zero-level set. Recent solutions [40, 77] propose a tailored weight function for UDF hierarchical sampling. Specifically, to upsample $N_{2}$ new points based on the original sampled points $\{q_{n}\}_{n=1}^{N_{1}}$, the positions of the upsampled points obey the posterior probability distribution,
$\begin{split}\Omega(n)&=(1-e^{-\tau_{n}\delta_{n}})\times\prod_{n^{\prime}=1}^{n-1}e^{-\tau_{n^{\prime}}\delta_{n^{\prime}}},\\ \tau_{n}&=\frac{s\cdot e^{-s\cdot f_{u}(q_{n})}}{(1+e^{-s\cdot f_{u}(q_{n})})^{2}},\end{split}$ (6)
where the sampling weight $\Omega_{n}$ is the cumulative probability over the interval $q_{n},q_{n+1}]$, that is, the probability that a newly generated upsampled point falls between $q_{n}$ and $q_{n+1}$. $\delta_{n}$ is the interval between $q_{n}$ and $q_{n+1}$, $\tau_{n}$ denotes the sampling density at $q_{n}$ and $s$ is a fixed standard deviation, respectively. We observe that this design introduces significant bias before the ray-surface intersection, which is primarily due to the lack of 3D awareness, similar as the analysis in Sec. III-B. As the example shown in Fig. 9 (a), due to the cumulative multiplicative effect, the high sampling densities before the ray-surface intersection result in significantly diminished sampling weights at the actual intersection. The bias of sampling weights consequently induces a concentrated distribution of upsampled points in the erroneous region before the ray-surface intersection.
Figure 11: Visualization of upsampled points based on our proposed point sampling prior and uniform upsampling strategy.
To mitigate the bias in hierarchical sampling, we introduce an auxiliary point sampling prior $f_{w_{p}}$ as a mask function of sampling densities. Specifically, the point sampling prior takes a window of UDFs as input and outputs the probability that the window contains a ray-surface intersection. The training of $f_{w_{p}}$ is similar as that of $f_{w}$, while the difference is that we use mask images as optimizing objectives here because we want $f_{w_{p}}$ to produce binary outputs. The mask value is then multiplied with $\{\tau_{n}\}$ for subsequent hierarchical sampling. The reason that we do not directly use the trained volume rendering prior for $\{\tau_{n}\}$ is that the training process of volume rendering prior itself also relies on hierarchical sampling. As demonstrated in Fig. 9 (b), the refined sampling densities $\{\tau_{n}\}$ through our point sampling prior $f_{w_{p}}$ effectively concentrates the upsampled points near the intersection.
Furthermore, we delve into the inverse transform sampling process and introduce a regularization term to promote the uniform distribution of upsampled points. Given uniformly distributed values $\{u_{n}\}_{n=1}^{N_{2}}$ on $[0,1]$, the positions of the $N_{2}$ upsampled points are determined by $F^{-1}_{\Omega}(u_{n})$, where $F^{-1}_{\Omega}$ is the inverse cumulative distribution function (CDF) of the sampling weights $\Omega$. Since $F^{-1}_{\Omega}$ is a non-smooth step function under the situation of discrete sampling, the distribution of the upsampled points $\{q^{\prime}_{n}\}_{n=1}^{N_{2}}$ along the ray is highly uneven, which results in a perception degeneration of the radiance fields, as shown in Fig. 10 (a). To facilitate the uniformity of the upsampling process, we first group the newly generated points that fall within the same interval of the original sampled point as $\{q^{\prime}_{n,1},q^{\prime}_{n,2},...q^{\prime}_{n,m}\}$, where $q^{\prime}_{n,j}$ represents the $j$-th point that fall within $q_{n},q_{n+1}]$. Then we re-order the $m$ points such that the interval of consecutive points $q^{\prime}_{n,j}$ and $q^{\prime}_{n,j+1}$ is $\frac{\delta_{n}}{m+1}$, where $\delta_{n}$ is the distance between $q_{n}$ and $q_{n+1}$, and $m$ is the number of points that fall within $q_{n},q_{n+1}]$, as illustrated in Fig. 10 (b). Through this way, we ensure the uniform distribution of upsampled points, thereby facilitating a comprehensive perception in 3D space during sampling.
We provide an example of sampling on a real garment in Fig. 11, where the ray starts from the bottom of the garment, passes through the waist, and eventually exits at the shoulder. Sampling bias causes VRPrior to erroneously concentrate the upsampled points around the waist of the garment. Guided by our point sampling prior, the upsampling process successfully circumvents the pseudo ray-surface intersection. Eventually, with our uniform sampling strategy, the upsampled points are densely and uniformly distributed near the ray-surface intersection.
Figure 12: Visualization of our strategy of refining Gaussian Splatting using pretrained volume rendering priors. The Gaussians are pushed toward the direction of increasing rendering weights under dense queries.
III-F Refining Gaussian Splatting with VRPrior
Recently, Gaussian Splatting has achieved remarkable progress in novel view synthesis [31, 71, 23], yet Gaussian-based surface reconstruction still struggles with discreteness, sparseness, and off-surface drift. Although several methods alleviate these issues by introducing regularization terms to align Gaussians with the surface [76, 39, 25, 10], they either cannot be integrated with neural implicit representations or incur the degeneration of geometric details. The reason is that these constraints are self-distilled from the Gaussians themselves and lack explicit knowledge of the guidance toward true surfaces. To address this limitation, we propose to employ our pretrained volume rendering priors as a general Gaussian surface prior that pushes the Gaussians toward the true surfaces predicted by the volume rendering prior, thereby refining the reconstructed geometry.
Specifically, as illustrated in Fig. 12, given a trained Gaussian Splatting model, we optimize an unsigned distance function on it, like GaussianUDF [39]. The UDF field provides a coarse estimation of the surface from the discrete and sparse Gaussian representations. We emit rays $V_{k}$ from the training views $\{I_{j}\}_{j=1}^{J}$, uniformly sample query points $\{q_{n}\}$ along each ray to obtain their UDF values and compute the rendering weights distribution $\{w_{n}\}$ around the surface. We do not use hierarchical sampling here in order to avoid fluctuations caused by random sampling intervals. We then select Gaussians $\{G_{t}\}$ that are both close to the ray and near the surface, interpolate the rendering weights at their projected positions on the ray, and compute the direction in which the rendering weight increases. A surface loss is applied to the Gaussians to move them along the direction of increasing rendering weights,
$\begin{gathered}\mathcal{L}_{surf}=\|\mu_{t}-\text{sg}(\mu_{t}+r_{k}*(d^{*}-d_{t}))\|,\\ d^{*}=d_{n^{*}},\ n^{*}=\arg\max_{n}w_{n}\end{gathered}$ (7)
where $\mu_{t}$ is the center of Gaussian $G_{t}$, $r_{k}$ is the direction of ray $V_{k}$, $d_{t}$ is the depth of the projection point from $G_{k}$ onto $V_{k}$, $d_{n}$ is the depth of $\{q_{n}\}$ and $\text{sg}(\cdot)$ is stop-gradient operation. Although the UDFs are not perfectly accurate, they provide reliable optimization directions for refining the Gaussian distribution. The Gaussian model and UDF model are jointly finetuned in a progressive manner, continuously improving each other and refining the geometry over time. Compared with existing methods that directly pull Gaussians to the UDF zero-level set [39, 76], our approach leverages dense queries around the zero-level set as strong geometric cues. It refines the surface in a conservative and gradual manner, effectively preventing error accumulation caused by ambiguous pulling directions.
IV Experiments
IV-A Experiment Settings
IV-A1 Data for Learning Priors
@@
We select one object from each one of “car”, “airplane”, “bottle” categories in ShapeNet dataset [6] and three objects from DeepFashion3D dataset [89] to form our training dataset for learning volume rendering priors. There is no overlap between the selected object and the testing objects. Our ablation studies demonstrate that these objects are sufficient to learn accurate volume rendering priors with good generalization capabilities across various shape categories. For each object, we first convert it into a normalized watertight mesh and then render 100 depth images with $\times600$ resolution from uniformly distributed viewpoints on a unit sphere. We utilize the volume rendering priors pre-trained on these six shapes to report our results.
IV-A2 Datasets for Evaluations
We evaluate our method on five datasets including DeepFashion3D (DF3D) [89], DTU [27], Replica [56], Insects and real-captured datasets. For DF3D dataset, we follow [41] and use the same 12 garments from different categories. For DTU dataset, we use the same 15 scenes that are widely used by previous studies. And we use all the 8 scenes in Replica dataset. To further evaluate the effectiveness of our method on reconstructing open structures, we select 10 insect objects from the Objaverse-XL dataset [16] to form the Insects dataset [87]. These objects feature detailed geometric elements as well as thin structures and open surfaces such as wings, which presents a challenge for multi-view reconstruction with UDFs. Similar to the process of constructing data for learning priors, we render 100 RGB images with $\times1024$ resolution from surrounding viewpoints for each object to form the training set. We also report results on real scans from NeUDF [40] and the ones shot by ourselves.
IV-A3 Baselines
We compare our method with the state-of-the-art methods which use different differentiable renderers to reconstruct open surfaces, including NeuralUDF [41], NeUDF [40], NeAT [46], 2S-UDF [17] and our previous work VRPrior [77]. We also compare our methods with state-of-the-art Gaussian Splatting-based reconstruction methods, including 2DGS [25], GOF [73], GaussianUDF [39], PGSR [10], GeoSVR [34] and QGS [81].
IV-A4 Metrics
For DTU dataset and DF3D dataset, we use Chamfer Distance (CD) as the metric. For Replica dataset, we report CD, Normal Consistency (N.C.) and F1-score following previous works [72, 3]. Moreover, we report the rendering errors in Tab. I using depth L1 distance, mask errors with cross entropy and L1 distance.
IV-A5 Implementation Details
Our volume rendering prior network consists of MLPs with 256 hidden units, where $f_{w_{1}}$ comprises 3 layers and $f_{w_{2}}$ has 6 layers with skip connections. The UDF function $f_{u}$ is an 8-layer MLP with skip connections and the color function $f_{c}$ is a 2-layer MLP with 256 hidden units. We apply Softplus as the activation function after the last layer of $f_{u}$ to preserve the output non-negative. The learning rate and decay scheduler of $f_{u}$ and $f_{c}$ are the same as previous works [41, 62]. We remove the normal vectors from the input of the color network because the normals are ambiguous at the zero-level set due to the non-differentiability of UDFs. For the experiments described in Sec. III-F, we train the Gaussian Splatting methods for 30k iterations. We begin training the UDFs from 9k iteration. During the final 10k iterations, we refine Gaussians by casting rays with a batch size of 2048, where 64 points are sampled along each ray.
@@ IV-A5 Implementation Details
TABLE I: Numerical comparisons in all ShapeNet categories. “Train” denotes the results evaluated on VRP training sets.
@@
[…]
TABLE II: Quantitative evaluations on DF3D[89], DTU[27] and Replica[56] datasets.
@@
2S-UDF[17] 1.99 0.90 0.34 0.63 0.16
VRPrior[77] 1.71 0.85 0.04 0.90 0.80
2DGS[25] 3.81 0.80 0.13 0.85 0.57
GOF[73] 2.49 0.74 0.09 0.80 0.58
GaussianUDF[39] 1.60 0.68 0.07 0.91 0.77
PGSR[10] 3.25 0.56 0.07 0.90 0.83
GeoSVR[34] 2.91 0.47 0.08 0.90 0.82
QGS[81] 4.46 0.54 0.06 0.91 0.74
Ours 1.59 0.66 0.03 0.92 0.85
Figure 13: Visual comparisons on open surface reconstructions with error maps on DeepFashion3D [89] dataset.
Figure 14: Visual comparisons of error maps on DTU [27] dataset.
Figure 15: Qualitative comparisons on Replica [56] dataset. Our method outperforms other methods on complex indoor scenes while other UDF-based methods struggle to recover complete and smooth surfaces.
Figure 16: Visual comparisons of reconstructed meshes on Insects dataset. Our method shows clear advantages in reconstructing thin structures such as the wings and the legs.
TABLE III: Numerical results on DTU with patch loss.
[…]
NeuralUDF[41] 1.05 1.10 1.07
Ours 0.72 0.89 0.80
NeuralUDF (w/ ft) 0.83 0.67 0.75
Ours (w/ ft) 0.73 0.59 0.66
IV-B Comparisons with the Latest Methods
IV-B1 Results on ShapeNet
Tab. I reports numerical comparisons in the experiment in Fig. 2 in terms of 3 metrics. We report the averages and variances over all 55 shapes. All of the objects are normalized to the range [-1, 1], and the values are magnified 100 times. We render depth images and mask images by forwarding ground truth unsigned distances at the same set of queries as other methods with the learned prior. For the learnable convergence parameters in the baseline methods, we fix them to the average values observed at convergence on DTU dataset. These values provide a reasonable reference for the choice of the factors. It should be noted that a larger factor does not correspond to higher accuracy. For instance, the factor in NeuralUDF determines the location where the UDF is flipped into an SDF. A very large value may cause this flip to occur before true surface, resulting in depth rendering errors. We calculate the L1-error and cross entropy error between predicted images and ground truth ones. We also measure the depth discrepancy between the sampled point with maximum rendering weight and the true surface along each ray, denoted as “Peak Diff-L1”. We achieve the best accuracy among all renderers for UDFs and SDFs. Although our volume rendering prior is unbiased since it is trained with ground truth UDFs and depths, certain inaccuracies still occur in real rendering due to random hierarchical sampling and the ill definition of the UDF field near the surfaces. This phenomenon is common in all volume rendering frameworks. Even NeuS [62], which employs a theoretically unbiased mapping, exhibits noticeable depth inaccuracies in practical scenarios, as shown in the “Peak Diff-L1” column of Tab. I. We also report the results of “NeuS-SDF” and our method on VRP training sets, which demonstrate the strong generalization ability of our approach.
@@ IV-B2 Results on DF3D
We first report evaluations on DF3D (CD $\times 10^{-3}$). Numerical comparison in Tab. II indicates that our learned prior produces the lowest CD errors among all handcrafted renderers. The visual comparison in Fig. 13 details our superiority on reconstruction with error maps. We see that our prior helps the network to recover not only smoother surfaces at most areas but also sharper edges at wrinkles.
TABLE IV: Reconstruction Comparison in terms of Chamfer Distance ($\times 10^{-3}$) on Insects datasets.
Methods S00 S01 S02 S03 S04 S05 S06 S07 S08 S09 Mean
NeuS[62] 3.22 4.61 3.63 5.93 5.94 4.75 2.84 3.79 4.32 6.90 4.59
NeuralUDF[41] 3.05 4.65 10.16 10.16 9.56 3.48 3.68 2.77 3.13 1.52 5.22
NeAT[46] 6.40 5.74 2.85 12.15 6.32 4.94 3.98 5.71 7.21 6.31 6.16
NeUDF[40] 6.24 4.86 9.81 8.69 5.69 8.12 12.81 6.23 4.49 7.04 7.40
2S-UDF[17] 12.16 4.83 7.15 9.49 7.40 4.89 7.71 3.20 7.98 4.21 6.90
VRPrior[77] 2.99 5.23 3.40 5.77 5.45 4.80 3.84 2.66 4.09 2.14 4.04
2DGS[25] 3.38 3.85 5.73 3.29 3.96 2.69 1.99 4.29 2.76 6.37 3.55
GOF[73] 3.51 3.47 5.98 3.49 4.00 2.22 1.74 4.48 2.31 5.92 3.71
GaussianUDF[39] 3.64 4.57 4.31 2.88 3.87 3.19 5.18 7.13 3.77 4.89 4.34
PGSR[10] 3.25 4.22 2.69 3.32 3.72 3.96 2.78 3.54 3.25 3.99 3.47
GeoSVR[34] 2.98 4.01 2.36 3.17 3.59 3.55 2.65 3.63 3.49 3.80 3.32
QGS[81] 3.82 3.40 3.17 2.60 4.11 4.53 2.08 4.16 3.60 4.22 3.57
Ours 2.48 4.40 2.31 4.47 4.26 3.71 3.34 2.19 3.04 1.76 3.19
Figure 19: Visual comparisons of real scans between VRPrior and our method.
IV-B3 Results on DTU
Tab. II reports our evaluation on DTU. Note that NeAT and 2S-UDF rely on additional mask supervision on DTU dataset. Our reconstructions produce low CD errors with our pre-trained prior. Although the shapes used to learn the prior are not related to any scenes in DTU, our prior comes from sets of queries in a local window on a ray, which are more general to unsigned distances along a ray in unobserved scenes. Although some recent 3DGS-based reconstruction methods achieve lower numerical results, they are tailored for closed surfaces and degrades severely in open and thin structures. Visual comparisons in Fig. 14 detail our reconstruction accuracy. Note that NeuralUDF uses additional patch loss [42] in DTU dataset to fine-tune the resulted meshes, which is not the primary contribution of the differentiable renderer. Hence, for fair comparison, we report the results of NeuralUDF without fine-tuning across all datasets. However, we also integrate our method with patch loss and report the comparisons in Fig. 17 and Tab. III, including our method and NeuralUDF with or without patch loss fine-tune. The visualization results show that our method, when fine-tuned using patch loss, still outperforms NeuralUDF under the same experimental conditions.
IV-B4 Results on Replica